### v1

@@ 3 Monte Carlo Simulations
In this section, we examine the finite-sample performance of the proposed MSS estimators using Monte Carlo simulations. The simulation design is detailed in subsection 3.1, and the results are presented in subsection 3.2.
@@ 3.1 The Simulation Setting
Here, (3.1) specifies a VAR of order one for the assets $R_{K,t}$ in the benchmark MSS, where $\mu\cdot\mathbf{1}_{K}$ and $\mathbf{A}$ denote the mean vector and the autoregressive coefficient matrix, respectively. The returns of possible redundant assets $R_{N,t}$ are connected to those in the benchmark MSS via (3.2), where $\mathbf{a}$ and $\mathbf{B}$ are parameters used to define the actual MSS in accordance with Lemma 1.
To incorporate GARCH effects, we define $\eta_{t}\equiv(\eta_{K,t}^{\top},\eta_{N,t}^{\top})^{\top}$, where the $i$th component of $\eta_{t}$, denoted as $\eta_{i,t}$ for $i=1,\ldots,K+N$, satisfies:
$\eta_{i,t}$ $=$ $d_{i,t}v_{i,t},$
$d_{i,t}$ $=$ $(0.1+0.1\eta_{i,t-1}^{2}+0.8d_{i,t-1}^{2})^{1/2},$
@@
Three MSS estimators are investigated in the simulation study: the infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, the one-step estimator $\overline{\mathrm{MSS}}_{p}$, and the two-step estimator $\widehat{\mathrm{MSS}}_{p}$, as defined in (2.13), (2.15) and (2.17), respectively. For these estimators, we set $p=0.05$, and their finite-sample performance is assessed using 10,000 Monte Carlo replications.
@@ 3.2 Simulation Results
Guided by theoretical insights, we first examine the empirical probabilities of the estimated MSS containing the true MSS, as reported in the second to fourth columns of Table 1. Consistent with the main theory established in Section 2, the simulation results indicate that both the one-step and two-step MSS estimators include the true MSS with high probability across all combinations of $K$ and $N$, provided the sample size is moderately large (e.g., $T\geq 180$). This probability approaches 1 for the two-step estimator when the sample size increases to 240, and for the one-step estimator when the sample size reaches approximately $300$.
Compared with the one-step estimator, the two-step MSS estimator exhibits a higher probability of containing the true MSS, especially when $N$ (the size of redundant set) is small. As discussed in Section 2, this is because the critical value employed by the two-step estimator is no larger than that of the one-step estimator, and their difference becomes more pronounced as $N$ decreases. The infeasible MSS estimator is the one which the two-step MSS estimator tries to mimic. As we can see from Table 1, their properties become identical for all combinations of $K$ and $N$ when the sample size becomes large. On the other hand, the one-step estimator is almost identical to the infeasible estimator when $K=1$, but it is more likely to miss assets in the true MSS when both $N$ and $T$ are small.
Next, we evaluate the ability of the estimated MSS to identify the true MSS, measured as the probability that the MSS estimator exactly matches the true MSS. This property is particularly important in practice, as it demonstrates the effectiveness of the MSS estimators in reducing the number of assets while maintaining the mean-variance efficient frontier. The results, presented in the fifth to seventh columns of Table 1, show that both the one-step and two-step MSS estimators accurately identify the true MSS with high probability when the sample size is sufficiently large, across all combinations of $K$ and $N$ (e.g., $T\geq 180$). When the sample size $T$ reaches $240$, the probabilities of the two-step MSS estimator coinciding with the true MSS exceed $0.9$ and converge to the nominal value of $0.95$ as $T=300$.
@@
One-step Est. Two-step Est. Infeasible Est. One-step Est. Two-step Est. Infeasible Est.
@@
$T=120$ 0.266 0.512 0.722 0.264 0.493 0.686
$T=180$ 0.819 0.931 0.967 0.815 0.887 0.917
$T=240$ 0.969 0.992 0.996 0.964 0.943 0.946
$T=300$ 0.997 0.999 1.000 0.992 0.945 0.946
  • Note: This table presents the simulation results for the MSS estimators described in Section 2. The second to fourth columns display the empirical probabilities of the estimated MSS containing the true MSS, denoted as $\mathrm{P}(\mathrm{MSS}\subset\widehat{\mathrm{MSS}})$, for the one-step, two-step, and infeasible MSS estimators. The fifth to seventh columns report the empirical probabilities of correctly identifying the true MSS, denoted as $\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}})$, for the one-step, two-step, and infeasible MSS estimators, respectively. The significance level $p$ is set at 0.05. All empirical probabilities are calculated based on 10,000 Monte Carlo replications.
In contrast, the one-step MSS estimator is more aggressive in reducing the size of the estimated set. Its probability of identifying the true MSS is lower than that of the two-step estimator when both $N$ and $T$ are small, but surpasses the two-step estimator’s probability as $T$ increases. This higher probability for the one-step MSS estimator at large $T$ arises due to its failure to maintain exact size control. The two-step MSS estimator can also achieve a probability of identifying the true MSS greater than 0.95 with a large sample size by specifying a smaller significance level $p$, such as $p=0.01$. However, the trade-off is that this adjustment increases the likelihood of missing assets from the true MSS when the sample size is small, resembling the properties of the one-step estimator.
Figure 1: Empirical Distribution of the Size of the Estimated MSS
Note: This figure illustrates the empirical distribution of the size of the estimated MSS for four different settings. The upper panel corresponds to the settings of $(K,N,T)=(1,7,120)$ and $(K,N,T)=(1,7,240)$, respectively, while the lower panel corresponds to the settings of $(K,N,T)=(7,1,120)$ and $(K,N,T)=(7,1,240)$, respectively. The significance level $p$ is set at 0.05. All empirical probabilities are calculated based on 10,000 Monte Carlo replications.
To further investigate the size of the estimated MSS, we present the empirical distributions of the sizes of the one-step and two-step MSS estimators in Figure 1 for $(K,N)=(1,7)$ and $(7,1)$ at $T=120$ and $240$. The results indicate that the empirical distribution of the estimated MSS size is nearly identical for both estimators when the size of the redundant set is large i.e., $N=7$. However, as shown in the lower-left panel of Figure 1, the one-step estimator may exclude assets belonging to the true MSS, leading to an underestimation of its size when both $N$ and $T$ are small. [^6] Additionally, we present the empirical distribution of the size of the intersection between the estimated MSS and the true MSS (i.e., $\cap\text{MSS}$ MSS) in Figure 4 in the Appendix. These results further demonstrate that the two-step test generally performs better when $N$ is small.
In summary, the simulation results in Table 1 and Figure 1 align with our theoretical findings: when the sample size is sufficiently large, the estimated MSS not only contains but also coincides with the true MSS with high probability. These results strongly support the practical effectiveness of the proposed MSS estimators, particularly the two-step estimator, in real-world applications.
@@ 4 An Empirical Application
In this empirical application, we further shed light on the interaction between individual stock momentum and factor momentum by estimating the $\mathrm{MSS}$ for various combinations of momentum strategies. We address the following two research questions. First, do the individual stock momentum factor and the factor momentum independently contribute to the mean-variance efficiency when analyzed alongside other well-established factors? Second, what role does factor momentum play in constructing the mean-variance efficient frontier when it coexists with the individual stock momentum factor? In contrast to Ehsani and Linnainmaa (2022), our analysis emphasizes the relative importance of various momentum strategies in shaping the mean-variance efficient frontier within the corresponding factor portfolios. [^8] For the sake of brevity, we focus on the estimation results of the two-step MSS test conducted at a significance level of 0.05. If a momentum strategy is more likely to be included in the $\mathrm{MSS}$, it should be regarded as a critical component in mean-variance analysis.
@@ 4.1 Data
Size SMB Banz (1981) 1964/07-2019/12 0.16% 3.01%
Value HML Rosenberg, Reid, and Lanstein (1985) 1964/07-2019/12 0.31% 2.88%
Profitability RMW Novy-Marx (2013) 1964/07-2019/12 0.27% 2.21%
Investment CMA Titman, Wei, and Xie (2004) 1964/07-2019/12 0.32% 2.00%
Accruals ACC Sloan (1996) 1964/07-2019/12 0.22% 1.93%
Betting against beta BAB Frazzini and Pedersen (2014) 1964/07-2019/12 0.88% 3.34%
Cash flow to price CFP Rosenberg, Reid, and Lanstein (1985) 1964/07-2019/12 0.30% 2.56%
Earnings to price ETP Basu (1983) 1964/07-2019/12 0.31% 2.63%
@@
Long-term reversals LTR De Bondt and Thaler (1985) 1964/07-2019/12 0.20% 2.57%
Net share issues NSI Loughran and Ritter (1995) 1964/07-2019/12 0.27% 2.42%
Quality minus junk QMJ Asness, Frazzini, and Pedersen (2019) 1964/07-2019/12 0.42% 2.30%
Residual variance RVA Ang, Hodrick, Xing, and Zhang (2006) 1964/07-2019/12 0.24% 5.04%
Short-term reversals STR Jegadeesh (1990) 1964/07-2019/12 0.49% 3.15%
@@ 4.2 Empirical Results
To distinguish the relative importance of individual stock momentum and factor momentum in shaping the efficient frontier, we next estimate the $\mathrm{MSS}$ for portfolios comprising one stock momentum strategy, one factor momentum strategy, and the Fama-French five factors. Table 4 reports the estimation results, highlighting three key findings. First, factor momentum is included in the estimated $\mathrm{MSS}$ for all ten cases, underscoring its critical role in constructing efficient frontiers. Second, although factor regressions suggest that individual stock momentum is largely explained by factor momentum (Ehsani and Linnainmaa, 2022), it nonetheless contributes to mean-variance efficient frontiers. Third, the market, size, profitability, and investment factors remain relevant in efficient frontiers when momentum in individual factors is present. However, the momentum in the first ten principal component factors (PCM) further subsumes the contribution of the size factor in efficient frontiers.
@@
Our estimation procedure provides more detailed insights into the relative importance of assets within the full set compared to the information presented in the table. To illustrate, Figure 2 displays plots of $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ along with their uniform confidence bands, for all momentum factors corresponding to the third column of Panel A in Table 5. The uniform confidence bands for $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ are centered at zero, with their upper bounds defined as $T^{-1/2}cv_{0.95,T}^{b}\hat{s}_{i,1}^{b}$ and $T^{-1/2}cv_{0.95,T}^{b}\hat{s}_{i,2}^{b}$, respectively. According to the estimation procedure, an asset is included in the $\mathrm{MSS}$ if, and only if, $|\hat{\alpha}_{i}|$ or $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$, or both exceed their respective confidence bands.
@@
Figure 2: Graphical Illustration of the Two-step MSS Estimation Procedure
@@
From Panel B of Table 5, we observe that the size of the estimated $\mathrm{MSS}$ is considerably smaller than the total number of factors, highlighting the effectiveness of the proposed estimation procedure in reducing the number of assets required to replicate the mean-variance efficient frontier. Additionally, the estimated $\mathrm{MSS}$ consistently includes momentum factors across all three cases. Notably, PCM subsumes FTM when both are present in the full set. Moreover, IAM and MOM consistently contribute to the efficient frontier, whereas SRM, which appears in the estimated $\mathrm{MSS}$ in several cases in Panel A, is excluded in Panel B due to the presence of new factors in the full set. Furthermore, the estimated $\mathrm{MSS}$ also includes excess market return (MKT), size (SMB), accruals (ACC), betting against beta (BAB), quality minus junk (QMJ), and short-term reversal (STR), underscoring their significance in mean-variance analysis.
Figure 3: Graphical Illustration of the Two-step MSS Estimation Procedure
@@
From Figure 3, we observe that PCM, along with MKT and STR, are included in the estimated $\mathrm{MSS}$ because both their $|\hat{\alpha}_{i}|$ and $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$ values are significantly different from zero. These factors emerge as the most important assets in the full set in terms of their contributions to mean-variance efficiency. Two factors, BAB and QMJ, are selected due to their $|\hat{\alpha}_{i}|$ values being significantly above zero, even though their $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$ values fall within the uniform confidence bands. The remaining assets in the estimated $\mathrm{MSS}$—MOM, IAM, SMB, ACC, and RVA—are included because their $|1^{\top}_{d-1}\hat{\beta}_{i}-1|$ values are significantly different from zero. Based on the statistic $M_{i,T}$, the ranking of factors in the estimated $\mathrm{MSS}$ is as follows: PCM, ACC, QMJ, SMB, MKT, IAM, MOM, STR, RVA, and BAB.
@@ Appendix A.1 Proofs of the Main Results
This section presents the conditions required to establish the main results of the paper, namely Lemma 1, Theorems 1 and 2, along with their proofs. We begin by outlining the key conditions.
@@
(i) the eigenvalues of $\Omega_{d,i}\equiv\lim_{T\rightarrow\infty}\mathrm{Var}(T^{-1/2}\sum_{t\leq T}e_{i,t})$are bounded away from zero; (ii) the bandwidth $\ell$ satisfies $\log(T)^{2}(\ell T^{-1/2+1/\tilde{r}}+\ell^{-1})=o(1)$ for some constant $\tilde{r}\in(4,r)$; (iii) $\gamma_{T}^{-1}T^{-1/8}=O(1)$.
Assumption 2 provides sufficient conditions to establish the validity of the bootstrap procedure for the estimation and inference of the MSS. Assumption 2(i) ensures that the asymptotic variance-covariance matrix of the estimator $\hat{\theta}_{i}$ is nonsingular for any $i\leq d$. Assumption 2(ii) imposes restrictions on the bandwidth $\ell$ used in the MBB procedure, specifically requiring that $\ell$ grows slower than $\log(T)^{-2}T^{1/2-1/\tilde{r}}$ but faster than $\log(T)^{2}$. Lastly, Assumption 2(iii) provides an upper bound on $\gamma_{T}$, which is satisfied by our recommendation, $\gamma_{T}=\min\{p/2,$ $0.05/\log(\log(T))\}$.
We now present the proofs of Lemma 1, Theorems 1 and 2. Throughout this section, $C$ denotes a positive constant that is independent of $i$, $t$ and $T$, and may vary from line to line.
@@
Combininig the results in (A.7) and (A.8) with Lemma A2, we have
@@
Since $\overline{cv}_{1-p,T}^{b}=O_{p}(1)$ by Lemma A4(i), from (A.10) we obtain:
$\min_{i\in\mathrm{MSS}}M_{i,T}-\overline{cv}_{1-p,T}^{b}$$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2}-O_{p}(1)$ (A.11)
@@
where the second inequality is by (2.16). From (A.11), it follows that:
$\mathrm{P}\left(\mathrm{MSS}\subset\overline{\mathrm{MSS}}_{p}\right)=\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>\overline{cv}_{1-p,T}^{b}\right)\geq\mathrm{P}\left(C^{-1}-o_{p}(1)>0\right)\geq 1-o(1),$ (A.12)
which shows the first claim of the theorem. [^11] Since the $\mathrm{MSS}$ is nonempty, (A.12) implies that $\overline{\mathrm{MSS}}_{p}$ is nonempty wpa1. Therefore, the finite sample adjustment in Step 7 of the algorithm in Subsection 2.2, when applied to construct $\overline{\mathrm{MSS}}_{p}$, is asymptotically negligible.
For any $p\in(0,1)$, let $\overline{cv}_{1-p}^{u}$ denote the ($1-p$)th quantile of $\max_{i\leq d}\tilde{M}_{i}$. Then by Lemma A5, there exists a positive sequence $\delta_{T}=o(1)$ such that
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)\geq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p-\delta_{T}}^{u}\right)-o(1).$ (A.13)
@@
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p,T}^{b}\right)\geq\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}\leq\overline{cv}_{1-p-\varepsilon}^{u}\right)-o(1),$ (A.14)

### v3

@@ 3 Monte Carlo Simulations
In this section, we examine the finite-sample performance of the proposed $\mathrm{MSS}$ estimator using Monte Carlo simulations. The simulation design is detailed in subsection 3.1, and the results are presented in subsection 3.2.
@@ 3.1 The Simulation Setting
Here, (3.1) specifies a VAR of order one for the assets $R_{K,t}$ in the benchmark MSS, where $\mu\cdot\mathbf{1}_{K}$ and $\mathbf{A}$ denote the mean vector and the autoregressive coefficient matrix, respectively. The returns of possible redundant assets $R_{N,t}$ are connected to those in the benchmark $\mathrm{MSS}$ via (3.2), where $\mathbf{a}$ and $\mathbf{B}$ are parameters used to define the actual MSS in accordance with Lemma 1. To incorporate GARCH effects, we define $\eta_{t}\equiv(\eta_{K,t}^{\top},\eta_{N,t}^{\top})^{\top}$, where the $i$th component of $\eta_{t}$, denoted as $\eta_{i,t}$ for $i=1,\ldots,K+N$, satisfies:
$\eta_{i,t}=d_{i,t}v_{i,t}\text{, \ where }d_{i,t}=(0.1+0.1\eta_{i,t-1}^{2}+0.8d_{i,t-1}^{2})^{1/2},$
@@
The infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$ and the proposed estimator $\widehat{\mathrm{MSS}}_{p}$, as defined in (2.11) and (2.13) respectively, are investigated in the simulation study. For both estimators, we set $p=0.05$, and evaluate their finite-sample performance using 10,000 Monte Carlo replications.
@@ 3.2 Simulation Results
Guided by theoretical insights, we first examine the empirical probabilities of the estimated $\mathrm{MSS}$ containing the true $\mathrm{MSS}$, as reported in the second and third columns of Table 1. Consistent with the main theory established in Section 2, the simulation results indicate that both the infeasible and proposed $\mathrm{MSS}$ estimators include the true $\mathrm{MSS}$ with high probability across all combinations of $K$ and $N$, provided the sample size is moderately large (e.g., $T\geq 180$), and this probability approaches 1 when the sample size increases to $240$.
Next, we assess the accuracy of the estimated $\mathrm{MSS}$ in identifying the true $\mathrm{MSS}$, measured as the probability that the $\mathrm{MSS}$ estimator exactly matches the true MSS. This property is particularly important in practice, as it demonstrates the effectiveness of the $\mathrm{MSS}$ estimators in reducing the number of assets while maintaining the mean-variance efficient frontier. The results, presented in the fourth and fifth columns of Table 1, show that the proposed MSS estimator $\widehat{\mathrm{MSS}}_{p}$ can identify the true $\mathrm{MSS}$ with high probability when the sample size is sufficiently large, across all combinations of $K$ and $N$ (e.g., $T\geq 180$). When the sample size $T$ reaches $240$, the probability of $\widehat{\mathrm{MSS}}_{p}$ coinciding with the true $\mathrm{MSS}$ exceeds $0.9$ and converges to the nominal value of $0.95$ as $T$increases to $300$.
Notably, the finite-sample performance of $\widehat{\mathrm{MSS}}_{p}$ closely matches that of the infeasible estimator it aims to mimic. Specifically, when the number of benchmark assets $K$ is small (e.g., $K\leq 3$), the empirical probabilities that $\widehat{\mathrm{MSS}}_{p}$ either contains or exactly matches the true $\mathrm{MSS}$ are nearly identical to those of the infeasible $\mathrm{MSS}$ estimator. Performance differences between these two estimators only arise when $K$ increases and the sample size $T$ is relatively small (e.g., $T=120$). However, as $T$ grows sufficiently large (e.g., $T\geq 240$), the proposed estimator $\widehat{\mathrm{MSS}}_{p}$ achieves performance comparable to the infeasible estimator. These findings confirm that $\widehat{\mathrm{MSS}}_{p}$ effectively restores the ideal performance of the infeasible estimator in finite samples.
In summary, the simulation results in Table 1 strongly align with our theoretical findings: when the sample size is sufficiently large, the estimated $\mathrm{MSS}$ not only contains but also coincides with the true $\mathrm{MSS}$ with high probability. These results strongly support the practical effectiveness of the proposed $\mathrm{MSS}$ estimator in real-world applications.
@@
Proposed Est. Infeasible Est. Proposed Est. Infeasible Est.
@@
$T=120$ 0.662 0.719 0.620 0.681
$T=180$ 0.953 0.965 0.898 0.915
$T=240$ 0.994 0.997 0.936 0.944
$T=300$ 0.999 0.999 0.947 0.950
  • Note: This table presents the simulation results for the MSS estimators described in Section 2. The second and third columns display the empirical probabilities of the estimated MSS containing the true MSS, denoted as $\mathrm{P}(\mathrm{MSS}\subset\widehat{\mathrm{MSS}})$, for the proposed and infeasible MSS estimators. The fourth and fifth columns report the empirical probabilities of correctly identifying the true MSS, denoted as $\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}})$, for the proposed and infeasible MSS estimators, respectively. The significance level $p$ is set at 0.05. All empirical probabilities are calculated based on 10,000 Monte Carlo replications.
@@ 4 An Empirical Application
In this empirical application, we further shed light on the interaction between individual stock momentum and factor momentum by estimating the $\mathrm{MSS}$ for various combinations of momentum strategies. We address the following two research questions. First, do the individual stock momentum factor and the factor momentum independently contribute to the mean-variance efficiency when analyzed alongside other well-established factors? Second, what role does factor momentum play in constructing the mean-variance efficient frontier when it coexists with the individual stock momentum factor? In contrast to Ehsani and Linnainmaa (2022), our analysis emphasizes the relative importance of various momentum strategies in shaping the mean-variance efficient frontier within the corresponding factor portfolios. If a momentum strategy is more likely to be included in the $\mathrm{MSS}$, it should be regarded as a critical component in mean-variance analysis.
@@ 4.1 Data
Size SMB Banz (1981) 1964/07-2019/12 0.24% 3.03%
Value HML Rosenberg, Reid, and Lanstein (1985) 1964/07-2019/12 0.29% 2.83%
Profitability RMW Novy-Marx (2013) 1964/07-2019/12 0.27% 2.17%
Investment CMA Titman, Wei, and Xie (2004) 1964/07-2019/12 0.27% 2.00%
Accruals ACC Sloan (1996) 1964/07-2019/12 0.22% 1.91%
Betting against beta BAB Frazzini and Pedersen (2014) 1964/07-2019/12 0.83% 3.27%
Cash flow to price CFP Rosenberg, Reid, and Lanstein (1985) 1964/07-2019/12 0.27% 2.51%
Earnings to price ETP Basu (1983) 1964/07-2019/12 0.29% 2.58%
@@
Long-term reversals LTR De Bondt and Thaler (1985) 1964/07-2019/12 0.21% 2.52%
Net share issues NSI Loughran and Ritter (1995) 1964/07-2019/12 0.22% 2.38%
Quality minus junk QMJ Asness, Frazzini, and Pedersen (2019) 1964/07-2019/12 0.39% 2.25%
Residual variance RVA Ang, Hodrick, Xing, and Zhang (2006) 1964/07-2019/12 0.12% 5.02%
Short-term reversals STR Jegadeesh (1990) 1964/07-2019/12 0.50% 3.09%
@@ 4.2 Empirical Results
To distinguish the relative importance of individual stock momentum and factor momentum in shaping the efficient frontier, we next estimate the $\mathrm{MSS}$ for portfolios comprising one stock momentum strategy, one factor momentum strategy, and the Fama-French five factors. Table 4 reports the estimation results, highlighting three key findings. First, factor momentum is included in the estimated $\mathrm{MSS}$ for all ten cases, underscoring its critical role in constructing efficient frontiers. Second, although factor regressions suggest that individual stock momentum is largely explained by factor momentum (Ehsani and Linnainmaa, 2022), it nonetheless contributes to mean-variance efficient frontiers. Third, the market, size, profitability, and investment factors remain relevant to efficient frontiers when either momentum in individual factors or momentum in the first ten principal component factors (PCM) is present, whereas the contribution of the value factor is limited.
@@
Our estimation procedure provides more detailed insights into the relative importance of assets within the full set compared to the information presented in the table. To illustrate, Figure 1 displays plots of $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ along with their uniform confidence bands, for all momentum factors corresponding to the third column of Panel A in Table 5. The uniform confidence bands for $|\hat{\alpha}_{i}|$ and $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$ start at zero, with their upper bounds defined as $T^{-1/2}cv_{0.95,T}^{b}\hat{s}_{i,1}^{b}$ and $T^{-1/2}cv_{0.95,T}^{b}\hat{s}_{i,2}^{b}$, respectively. According to the estimation procedure, an asset is included in the $\mathrm{MSS}$ if, and only if, $|\hat{\alpha}_{i}|$ or $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$, or both exceed their respective confidence bands.
@@
Figure 1: Graphical Illustration of the MSS Estimation Procedure
@@
From Panel B of Table 5, we observe that the size of the estimated $\mathrm{MSS}$ is considerably smaller than the total number of factors, highlighting the effectiveness of the proposed estimation procedure in reducing the number of assets required to replicate the mean-variance efficient frontier. Additionally, the estimated $\mathrm{MSS}$ consistently includes momentum factors across all three cases. Notably, PCM subsumes FTM when both are present in the full set. Moreover, MOM and IAM consistently contribute to the efficient frontier, whereas SRM, which appears in the estimated $\mathrm{MSS}$ in several cases in Panel A, is excluded in Panel B due to the presence of new factors in the full set. Furthermore, the estimated $\mathrm{MSS}$ also consistently includes excess market return (MKT), size (SMB), accruals (ACC), quality minus junk (QMJ), and short-term reversal (STR), underscoring their significance in mean-variance analysis.
Figure 2: Graphical Illustration of the MSS Estimation Procedure
@@
From Figure 2, we observe that MKT and STR are included in the estimated $\mathrm{MSS}$ because both their $|\hat{\alpha}i|$ values and their $|1^{\top}{d-1}\hat{\beta}i-1|$ values are significantly different from zero. In contrast, PCM, MOM, IAM, SMB, ACC, and RVA are included solely because their $|1^{\top}{d-1}\hat{\beta}_{i}-1|$ values are significantly different from zero, even though their $|\hat{\alpha}_{i}|$ values fall within the uniform confidence bands. Additionally, QMJ is selected exclusively due to its $|\hat{\alpha}i|$ value being significantly above zero. Based on the statistic $M{i,T}$, the factors included in the estimated $\mathrm{MSS}$ are ranked as follows: PCM, ACC, QMJ, SMB, MKT, IAM, MOM, STR, and RVA. This ranking further highlights PCM’s critical contribution to mean-variance efficiency.
@@ Appendix A.1 Proofs of the Main Results
This section presents the conditions required to establish the main results of the paper, namely Lemma 1 and Theorem 1, along with their proofs. We begin by outlining the key conditions.
@@
(i) the eigenvalues of $\Omega_{d,i}\equiv\lim_{T\rightarrow\infty}\mathrm{Var}(T^{-1/2}\sum_{t\leq T}e_{i,t})$are bounded away from zero; (ii) $\log(T)^{2}(\ell T^{-1/2+1/\tilde{r}}+\ell^{-1})=o(1)$ for some constant $\tilde{r}\in(4,r)$.
Assumption 2 provides sufficient conditions to establish the validity of the bootstrap procedure for the estimation and inference of the MSS. Assumption 2(i) ensures that the asymptotic variance-covariance matrix of the estimator $\hat{\theta}_{i}$ is nonsingular for any $i\leq d$. Assumption 2(ii) imposes restrictions on the bandwidth $\ell$ used in the MBB procedure, specifically requiring that $\ell$ grows slower than $\log(T)^{-2}T^{1/2-1/\tilde{r}}$ but faster than $\log(T)^{2}$.
We now present the proofs of Lemma 1 and Theorem 1. Throughout this section, $C$ denotes a positive constant that is independent of $i$, $t$ and $T$, and may vary from line to line.
@@
Combining the results in (A.7) and (A.8) with Lemma A2, we have
@@
Let $cv_{1-p,T}^{b}\equiv cv_{1-p,T}^{b}(\mathcal{M}\backslash\widehat{\mathrm{MSS}}_{p})$. Since $cv_{1-p,T}^{b}(\mathcal{M})=O_{p}(1)$ by Lemma A4 and $0\leq cv_{1-p,T}^{b}\leq cv_{1-p,T}^{b}(\mathcal{M})$, it follows from (A.10) that
$\min_{i\in\mathrm{MSS}}M_{i,T}-cv_{1-p,T}^{b}$$\geq$ $(C^{-1}-o_{p}(1))T^{1/2}\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)^{1/2}-O_{p}(1)$ (A.11)
@@
where the second inequality is by (2.15). From (A.11), we obtain:
$\mathrm{P}\left(\mathrm{MSS}\subseteq\widehat{\mathrm{MSS}}_{p}\right)=\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p,T}^{b}\right)\geq\mathrm{P}\left(C^{-1}-o_{p}(1)>0\right)\geq 1-o(1),$ (A.12)
which shows the first claim of the theorem. [^10] Since the $\mathrm{MSS}$ is nonempty, (A.12) implies that $\widehat{\mathrm{MSS}}_{p}$ is nonempty wpa1. Consequently, the finite sample adjustment described in the paragraph following the algorithm in Subsection 2.2 is asymptotically negligible.
To establish the second claim of the theorem, we first observe that since $\mathcal{S}\subseteq\mathcal{S}^{\prime}$ implies $\max_{i\in\mathcal{S}}M_{i,T}^{b}\leq\max_{i\in\mathcal{S}^{\prime}}M_{i,T}^{b}$, it follows that
$cv_{1-p,T}^{b}(\mathcal{S})\leq cv_{1-p,T}^{b}(\mathcal{S}^{\prime}).$ (A.13)
From this and (A.12), we have $cv_{1-p,T}^{b}\leq cv_{1-p,T}^{b}(\mathrm{MSS}^{c})$ wpa1, where $\MSS\mathrm{MSS}^{c}\equiv\mathcal{M}\backslash\mathrm{MSS}$. Thus
$\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}\right)\leq\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS}^{c})\right)+o(1).$ (A.14)
By Lemma A5, there exists a positive sequence $\delta_{T}=o(1)$ such that
$\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS}^{c})\right)\leq\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p+\delta_{T}}^{u}(\mathrm{MSS}^{c})\right)+o(1).$ (A.15)
@@
$\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p,T}^{b}(\mathrm{MSS}^{c})\right)\leq\mathrm{P}\left(\max_{i\in\mathrm{MSS}^{c}}M_{i,T}\leq cv_{1-p+\varepsilon}^{u}(\mathrm{MSS}^{c})\right)+o(1),$ (A.16)